This one is bogging me down. I need help.
Driving marathon. Felix drove 800 miles in x hours on Monday.
a) write a rational expression for his average speed.
b)On Tuesday, he drove for six hours at the same average speed. Write a rational expression for his distance on Tuesday.

Answer by stanbon(48558)

General formula: distance = rate*time
--------------------------------------------

Driving marathon. Felix drove 800 miles in x hours on Monday.
a) write a rational expression for his average speed.
rate = distance/time = 800/x mph
--------------------------------------
b)On Tuesday, he drove for six hours at the same average speed. Write a rational expression for his distance on Tuesday.
800/6 = 400/3 mph
==========================
Cheers,
Stan H.